OpenStudy (anonymous):

michael worked 16 hours last week. he earned $5 per hour at a restaurant and $5.50 per hour at a grocery store. if he earned a total of $82, how many hours did he work at the grocery store?

OpenStudy (anonymous):

\[$ {5 x + 5.5(16-x)} = $ 82\] \[5x + 88 - 5,5x = 82\] \[-0.5 x = -6\] \[x = \frac{6}{0.5} = 12 hrs\] He worked 12 hours at Resturant and 4 hours at grocery store